OpenStudy (anonymous):

write sin(x)+2cos(x) in terms of sine only

OpenStudy (anonymous):

Take a look at the graphs of sin(x) and cos(x): http://puu.sh/8nz0m.png You should notice that cos(x) is just sin(x) shifted to the left pi/2 units. So, we want to shift cos(x) to the right pi/2 units. Since we're shifting to the right, we need to account for this inside the argument of cos(x). cos(x-pi/2) = sin(x). That's the relationship between these two functions. But we're writing cos(x) in terms of sin(x). So, we should shift sin(x) to the left pi/2 units. Here is the other relationship: sin(x+pi/2) = cos(x). So now, sin(x) + 2cos(x) = sin(x) + 2(sin(x+pi/2)).